OpenStudy (anonymous):

What is the probability of obtaining exactly 4 successes in 5 trials when the probability of success is 0.5. Round your answer to 4 decimal places.

OpenStudy (anonymous):

i'm not so sure but i think the probability of success is 0.5, then the probability of not success is 1-0.5=0.5 obtaining 4 successes in 5 trials = 4 successes and 1 not success so, the probability of it is 0.5^4 x 0.5 = 0.5^5 = 1/32

OpenStudy (anonymous):

1/32 = 0.0312

OpenStudy (gg):

I think you are right. Those happenings are independent, so we just multiply probabilities of each.

OpenStudy (anonymous):

thats actually not the answer its (5)(0.5)^4(0.5)^5-4=0.1563

OpenStudy (gg):

5 times (0.5^4) times (0.5^5) -4 ?

OpenStudy (anonymous):

hmm maybe because there are 5 trials, the 1/32 that i got before should be multiplied by 5, therefore the answer is the same as yours ^_^

OpenStudy (anonymous):

1/32 x 5 = 0.03125 x 5 = 0.15625

OpenStudy (anonymous):

yes

OpenStudy (gg):

but why to multiply it by 5?

OpenStudy (anonymous):

i'm also confused, but i think it's because there are 5 trials

OpenStudy (gg):

but if we have 100 trials, than the possibility would be 3.1, and that is impossible, right? so, that is not the reason :)

OpenStudy (anonymous):

i figured it out by the book saying this Let T denote the outcome “Tail shows” (success) and let H denote the outcome “Head shows” (failure). Using Formula (2), in which k = 1 (one success), n = 6 (the number of trials), and (the probability of success using a fair coin), we obtain

OpenStudy (gg):

P= (n!/[k!(n-k)!] ) p^k * q^(n-k) use this formula. n is number of trials, k is number of successes, p is possibility of success, q=1-p :) is it ok now? :)

OpenStudy (anonymous):

oh yes, combination. 5C4 0.5^4 * 0.5. thank you

OpenStudy (gg):

problem solved :)